Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
nardecchia, pinto, gouaty, magazzu, tringali - 22:51 Friday 22 September 2023 (61755) Print this report
Comment to DAS tuning (61754)

Today afternoon, it was not possible to reach LN2 most probably due to the unbalance of the B1p 56 MHz upper and lower sidebands (see fig.1).

In the lock started at 18.11 UTC (when CARM NULL 1F has been achieved), we added -11 as MICH offset in order to overimpose the sidebands  to proceed up to LN2, the trick worked (see fig.2) but the ITF unlocked at LN2 due to huge oscillation at ~55 Hz .  

Thus, to save the weekend, we decided to reverse the last step of DAS WP.

So, at 19.24 UTC, we came back to STEP_1S (61729) .

 

STEP_1S
19.24 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.090

1.857

on the pickoff

0.418

 0.0146

0.302

N

on the ITF

0.091

0.266 

2.972

on the pickoff

0.561

0.044

0.484

After that, the ITF achieved LN3 (at 20.25 UTC) without adding extra offsets (fig.3).

As requested by Romain, the BS TY SET was put to 0.

We left the ITF to Romain.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×